OpenStudy (anonymous):

Solve for x: log2(x+1) + log2(x+3) = 3 I got x = -5, 1... I just need someone to check this :)

OpenStudy (anonymous):

8=(x+1)/(x+3) ===? x+1=8x+24 ===> 7x=-23 ===> x= -23/7

OpenStudy (lalaly):

\[\log_2 {(x+1)(x+3)} = 3\]\[2^3 = (x+1)(x+3)\]\[x^2+4x+3=8\]\[x^2+4x-5=0 \]\[(x+5)(x-1)=0\]\[x=-\]

OpenStudy (lalaly):

x=-5, x=1

OpenStudy (lalaly):

yu are right

OpenStudy (anonymous):

Thanks everyone :)

OpenStudy (lalaly):

=) yw